OpenStudy (anonymous):

What are the x– and y–intercepts of y = 2x2 – 10x + 8? Can someone show me how to get the answer?

OpenStudy (savvy):

x int. = 3 y int. = 8

OpenStudy (anonymous):

To find intercept with the y-axis you set x to 0 (because the y-axis is located where x=0) and solve for y. y = 2(0)^2 - 10(0) + 8 = 8, so intercept is at (0,8) To find intercept with the x-axis you set y to 0 (because the x-axis is located where y=0) and solve for x. 2x^2 - 10x + 8 = 0 (2x - 2)(x - 4) = 0 2x -2 = 0 or x - 4 = 0 so intercepts at (1,0) and (4,0)